The demand for sausages is Q = 280 − 8P and the supply of sausages is Q = 12P + 20. A per-sausage tax of $2 is levied on vendors. Calculate the after-tax price paid by consumers and the after-tax price received by sausage vendors. What would the price be without the tax?
Answer #1

Answer)

Without tax price, D=S

280-8P = 12P+20

20P = 260

P=13(Without Tax Price)

Qd= 280-8(13) =176

Qs= 12(13) +20 = 176

After tax price,

Q= 280-8(P+2)

176 = 280-8P-16

8P=88

P=11(Price of buyers)

Q=12(P-2) +20

176=12P-24+20

12P = 180

P = 15(Price of sellers)

Tax is levied on vendors so producers have to bear greater prices and consumers will have to bear low prices.

If you have any doubts please comment...